Fermi National Laboratory
 

Operation Summary for June 23, 2010 - June 25, 2010

 

Wednesday June 23

The day shift began with Operations monitoring store 7911 (89.59E30) and the Recycler stash (291.1E10), with the Antiproton Source (Pbar) stacking (11.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson FTTB T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7911 at 9:45 AM. While the TeV was at 150 GeV, technicians worked on a TeV sector B3 cryo system valve (B3EVLH) and the sector D1 BPMs. TeV personnel conducted antiproton-only store studies. Afterwards, cryo system techs began replacing the TeV sector F1 wet engine flywheel. Operations reported several NuMI beam permit trips due to the NuMI extraction kicker temperature; NuMI personnel investigated. MI personnel conducted collimator studies.

During the evening shift, operators began the TeV cryo system recovery at F1. Operations disabled the T-977 beam at 6 PM. Operations reported that a line of nasty storms blew through and caused glitches that tripped various machines. The TeV F3 wet engine controller failed during the storm (probably lightning induced) and needed cryo system personnel to get it running again. Pbar suffered a storm-induced problem with a Pbar Debuncher RF station (DRF1) and the laser light link in the tunnel; Pbar personnel accessed and repaired the laser. Operations reported that the F1 cryo system recovered and operators put the TeV through a dry squeeze at 9:58 PM. The TeV C3 sump pump alarmed on a high water indication that did not clear within an hour; operators accessed and found the sump pit empty.

Thursday June 24

The midnight shift began with Operations monitoring the stash (494.0E10), with Pbar stacking (21.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators continued the TeV check out and turn-on and put the TeV through a dry squeeze at 1:45 AM. Operators checked the orbits and discovered a significant vertical orbit distortion at flattop caused by a power supply (VD18). They restored the appropriate settings and had the TeV ready for a HEP shot setup by 5 AM. Operations established store 7918 at 6:45 AM with an average initial luminosity (AIL) of 293E30. Operations contacted Pbar personnel to investigate the lower than normal stacking rates, presumably due to changes in the 4-8 core cooling system.

During the day shift, Operations terminated store 7918 at 3:04 PM. Operators prepared the TeV A-E enclosures and Switchyard Enclosure B for access to inspect and repair the vacuum at D2 and investigate a high water alarm. D0 and CDF also accessed their collision halls. RF personnel swapped out the modulator for MIRF1 and operators returned the station to service. MIRF18 faulted off; RF personnel replaced a faulty cavity controller.

During the evening shift, accesses into CDF, D0, Enclosure B, and TeV A-E continued. Operations disabled beam to T-977. Operations reported that the Pelletron developed a problem with its solenoid power supply. Electron Cooling system personnel reconfigured the system in order to avoid making a tank access. All accesses came to an end and Operations had the TeV ramping towards a dry squeeze by 8:15 PM. Operations put the TeV into shot setup and established store 7922 at 11:40 PM with an AIL of 328E30.

Friday June 25

The midnight shift began with Operations monitoring store 7922 (279.44E30) and the stash (178.5E10), with Pbar stacking (9.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that despite a few Pelletron trips, the system remained stable since its reconfiguration. Operations established beam to T-977.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/25/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory